OpenStudy (anonymous):

50th derivative of y=cos2x. Ive seen examples but i still dont understand why 2 is being raised to a power instead of being multiplied

OpenStudy (anonymous):

it is being multiplied. it is being multiplied by itself

OpenStudy (anonymous):

i found this on a different website and they posted this f^1(x) = -1(2)^1sin(2x) f^2(x) = -1(2)^2 cos(2x) f^3(x) = 1(2)^3sin(2x) f^4(x) = 1(2)^4cos(2x) says its being raised to it? isnt the answer for the 50h derivative for y=cos2x; y=-2^50cos2x

OpenStudy (anonymous):

\[-2\sin(2x)\]\[-4\cos(2x)\]\[8\sin(2x)\]\[16\cos(2x)\]...

OpenStudy (anonymous):

i still dont get it